Sign in
Please select an account to continue using cracku.in
↓ →
In a defective 6-faced die with numbers 1 to 6 inscribed, the probability of getting an odd number is twice the probability of getting an even number. Find the probability of getting a 2-digit prime number on 2 successive throws of the die.
If 'x' is the probability of rolling 2/4/6, then '2x' will be the probability of rolling any odd number
=> x + x + x + 2x + 2x + 2x = 1 => x = 1/9
Two-digit prime numbers that can be formed with 1, 2, 3, 4, 5, 6 are:
11, 13, 23, 31, 41, 53, 61
=> Required probability = 2/9*2/9 + 2/9*2/9 + 1/9*2/9 + 1/9*2/9 + 2/9*2/9 + 1/9*2/9 = 18/81 = 2/9 [No correct option]
Was this solution helpful?
Book Free CAT Mentorship
Get personalized CAT strategy from a 99%iler
500+ students mentored
OTP Verification
Enter the 6-digit code sent to your phone
Booking Summary
Enter OTP
Didn't receive the OTP?
Start your IIM journey with the right preparation and crack CAT 2026.
Educational materials for CAT preparation